OpenStudy (anonymous):

Write g(x) = 3cos(5x) +3sqroot3sin(5x) in terms of sine only. Please help.

OpenStudy (anonymous):

\[b\cos(x)+a\sin(x)=\sqrt{a^2+b^2}\sin(x+\theta)\] for the appropriate \(\theta\)

OpenStudy (anonymous):

find \(\theta\) via \[\sin(\theta)=\frac{a}{\sqrt{a^2+c^2}}\] and \[\cos(\theta)=\frac{b}{\sqrt{a^2+b^2}}\]

OpenStudy (anonymous):

I get the answer 6sin(5x+pi/3) but, its wrong

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

yeah i screwed up that is why

OpenStudy (anonymous):

it is \[\cos(\theta)=\frac{a}{\sqrt{a^2+b^2}}\] and \[\sin(\theta)=\frac{b}{\sqrt{a^2+b^2}}\]

OpenStudy (anonymous):

this gives \(\theta=\frac{\pi}{6}\) i had it backwards

OpenStudy (anonymous):

thank you (: